on a test of 80 items, pedro got 93% correct. (there was partial credit on some items.) how many items did he get correct?incorrect?
Answer by oberobic(2304)

The percent correct is the number correct divided by the total number of items.
If there were 100 questions, then 93/100 would yield 93% correct.
But instead of 100 we have only 80 questions.
Since we want the number correct, 'x', to be 93% of 80, we can setup the equation as:
x/80 = 93/100
Multiply both sides by 80
x = 80(93/100) = (80*93)/100 = 7440/100 = 74.40
So Pedro got 74 and 4/10 questions correct.
...
Another way to get there is to just multiply 80 * .93 = 74.40